Jetson Spacecraft Corp. shows the following information on its 2015 income statement: sales = $300888; costs = $218637; other expenses = $8585; depreciation expense = $18372; interest expense = $14158; taxes = $17217; dividends = $11001. In addition, you’re told that the firm issued $6036 in new equity during 2015 and redeemed $4781 in outstanding long-term debt. If net fixed assets increased by $22082 during the year, what was the addition to NWC?
Answer #1

Operating cash flow = EBIT + Depreciation - Taxes = (Sales- Costs-other expenses - depreciation) - taxes + depreciation
= (300888 -218637 -8585 -18372) -17217 + 18372 = 56449

Cash Flow from Assets = Cash flow to debtors + Cash flow to equity holders = (Interest - New Debt) + ( dividends- new Equity )
= (14158 - 4781) + (11001 - 6036 ) = 14,342

Cash flow from Assets = OCF - Increase in Fixed = Increase in NWC
14,342 = 56449 - 22082 - Increase in NWC
Increase in NWC = 20,025

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well